# KnowPhish: Large Language Models Meet Multimodal Knowledge Graphs for Enhancing Reference-Based Phishing Detection

### v1

@@ Abstract
Phishing attacks have inflicted substantial losses on individuals and businesses alike, necessitating the development of robust and efficient automated phishing detection approaches. Reference-based phishing detectors (RBPDs), which compare the logos on a target webpage to a known set of logos, have emerged as the state-of-the-art approach. However, a major limitation of existing RBPDs is that they rely on a manually constructed brand knowledge base, making it infeasible to scale to a large number of brands, which results in false negative errors due to the insufficient brand coverage of the knowledge base. To address this issue, we propose an automated knowledge collection pipeline, using which we collect and release a large-scale multimodal brand knowledge base, KnowPhish, containing 20k brands with rich information about each brand. KnowPhish can be used to boost the performance of existing RBPDs in a plug-and-play manner. A second limitation of existing RBPDs is that they solely rely on the image modality, ignoring useful textual information present in the webpage HTML. To utilize this textual information, we propose a Large Language Model (LLM)-based approach to extract brand information of webpages from text. Our resulting multimodal phishing detection approach, KnowPhish Detector (KPD), can detect phishing webpages with or without logos. We evaluate KnowPhish and KPD on a manually validated dataset, and on a field study under Singapore’s local context, showing substantial improvements in effectiveness and efficiency compared to state-of-the-art baselines.  [^1] Preprint. Under review. Codes and data will be released at a later date.
@@ 1 Introduction
Phishing attacks are one of the most impactful type of scams, harming both individuals and businesses: in 2023, an estimated \$1.026 trillion was lost by consumers worldwide in scams [5], of which phishing scams are among the most common types. Phishing scams also account for about 90% of data breaches for organizations [11]. This problem has been exacerbated by the proliferation of automated phishing kits, enabling malicious actors to easily mimic genuine pages while evading detection using countermeasures [64]. Accordingly, the number of phishing attacks has increased by 47.2% in 2022 compared to the previous year [65]. These underscore the importance and urgency of tackling the issue, and the need for effective automatic phishing detection approaches.
@@
Many efforts have been made to counter phishing attacks, including anti-phishing blacklists, heuristic-based, and feature-based models [14]. Blacklist-based methods [40, 42, 45] compare an input URL with a predefined blacklist of malicious URLs, but these are reactive approaches that require phishing sites to be first reported or detected through other means. Heuristic-based [13] and feature-based models [28, 37, 18, 61, 31, 54, 35] extract features to proactively identify new phishing webpages. However, as these do not utilize logo information, they are greatly limited in their ability to detect plentiful phishing pages which are mainly identifiable by the presence of a logo [33]. Moreover, as they depend on statistical features for detecting phishing pages, they are susceptible to distribution shift due to the constantly changing nature of phishing campaigns.
@@
In contrast, reference-based phishing detectors (RBPDs), which work by comparing the logos on a target webpage to a known set of logos, have been established as the state-of-the-art phishing detection paradigm, garnering considerable research attention [1, 32, 33]. Specifically, a RBPD consists of brand knowledge base (BKB) containing brand information (the logos and legitimate domains of brands) and a detector backbone which uses information from this BKB for phishing detection. To detect if a webpage is a phishing or benign page, RBPDs first identify the webpage’s brand intention, i.e., the brand that the webpage presents itself as (e.g., a webpage with a Adobe logo and appearance has the brand intention of Adobe). Then, if the webpage is detected to have an intention of a certain brand, but its domain does not match the legitimate domains of that brand, the webpage is classified as a phishing. As virtually all phishing pages do not utilize the legitimate domains of the original brand, this approach typically obtains high precision. Moreover, as this approach is based on an invariant that does not change over time, it is relatively robust to distribution shift [32, 33].
Challenges.
Figure 1: A logo-less phishing website showing its brand intention through text, visible in the page’s HTML.
@@
Despite the advantages of RBPDs, they have two major limitations which we focus on. (1) The first limitation is Limited-Scale BKB: RBPDs fundamentally rely on their BKB to identify the brand intention of a website. However, it is labor-intensive to construct and maintain a large-scale BKB manually. Phishpedia [32] and PhishIntention [33] rely on manual curation, hence are limited to a small BKB of 277 brands. A recent method, DynaPhish [34], proposes to dynamically expand the BKB during deployment time. However, this leads to extremely long running time, e.g., averaging $10.6$ seconds per sample in our experiments. Moreover, it cannot recognize logo variants of those unseen brands, leading to potential false negatives. (2) The second limitation is Textual Brand Intention: Phishing webpages can convey their brand intention via text, instead of via logos. Figure 1 shows a webpage without a logo, but its target brand, Australia Post, is presented in the HTML. Its brand intention is clear to visitors of the page, but existing RBPDs cannot identify it because they solely operate within the image modality.
Present Work.
@@
In this paper, we seek to address both of these issues. First, we propose an automated knowledge collection pipeline, with which we construct and publicly release a large-scale multimodal BKB named KnowPhish containing rich logo variants, aliases, and domain information of each brand. KnowPhish is constructed based on our empirical analysis which finds that phishing targets mostly belong to a few high-value industries, and this is stable over time. Hence, using brand-industry relations modelled from a publicly available knowledge base, Wikidata, we search for a set of potential phishing targets and their brand knowledge predictively, leading to a BKB covering around 20k potential phishing targets. KnowPhish can be used to boost the performance of existing RBPDs in a plug-and-play manner, as demonstrated in our experiments.
@@
Next, to address the issue of textual brand intention, we develop a Large Language Model (LLM)-based approach to identify the brand intention of webpages in conjunction with the alias information in our KnowPhish KB. Our approach can be directly integrated with any standard visual RBPD [33], augmenting it to form a multimodal RBPD which detects brand intention through both visual and textual modalities. Our resulting multimodal phishing detection approach, named KnowPhish Detector (KPD), is able to detect phishing webpages with or without logos.
@@
We then evaluate the effectiveness and efficiency of KnowPhish and KPD on TR-OP, a manually validated dataset that comprises 5k benign and 5k phishing webpages, which we release publicly. We also evaluate our approach on a field study under Singapore’s local context, to study how well different approaches generalize to such a local context. The resulting data, which we call SG-SCAN, contains 10k webpages from Singapore local webpage traffic over 6 months. On experiments on the two datasets, KnowPhish significantly boosts the effectiveness of RBPDs, and is 30 or more times faster than the on-deployment framework DynaPhish [34], when equipped with image-based RBPDs. Moreover, incorporating our multimodal approach, KPD, can substantially boost the number of detected phishing webpages.
@@
  • Multimodal Brand Knowledge Base. We propose KnowPhish, the first-of-its-kind large-scale multimodal BKB for phishing detection, and its automated construction approach. KnowPhish can be used in any RBPDs to boost their brand knowledge in a plug-and-play manner.
  • Multimodal Reference-based Phishing Detector. We propose an LLM-based approach to identify textual brands from HTML to handle logo-less phishing webpages. Our approach directly integrates with any existing RBPD, augmenting it to form a multimodal RBPD which can detect logo-less phishing webpages.
  • Effectiveness and Efficiency. Extensive experiments show that our approach significantly enhances effectiveness and efficiency of phishing detection compared to state-of-the-art baselines. We also demonstrate its effectiveness on a field study, and validate the robustness to adversarial attacks.
@@ 2 Formalization
Threat Model.
In a phishing attack, a malicious actor misleads visitors into believing that their webpage comes from a legitimate brand, misleading users into providing their credentials (such as username and password). Note that we do not consider other types of attacks, such as malware or miscellaneous scams that do not fit the above description, to be within our scope.
@@
Reference-Based Phishing Detection.
Since phishing webpages need to convey brand intention, the state-of-the-art RBPD approach relies on identifying this brand intention, by comparing images on the page to a set of known reference logos. Formally, an RBPD consists of a brand knowledge base (BKB) and a detector backbone utilizing this BKB.
@@
[…]
Evasion Attacks.
The attackers may attempt to bypass $f_{\mathcal{B}}$ via the following methods:
T1: Phishing with Logo Variants. To circumvent the online knowledge expansion approach for $\mathcal{B}$ (e.g., DynaPhish[34]), attackers can use other legitimate logo variants of $b$ instead of the one displayed on $b$’s official webpage.
T2: Phishing with Text Brands. Instead of using a logo to present their brand intention, the attacker can rely on text w.html to show its brand intention, making image-based phishing detectors completely fail.
T3: HTML Obfuscation. Phishing attackers may employ evasion techniques on HTML content to hinder effective information extraction by text-based methods.
We address T1 by constructing a large-scale multimodal BKB with rich logo information, i.e., multiple logos per brand. We address T2 by developing an LLM-based approach to extract text brands from HTML. For T3, we show in Section 5.5 that our LLM-based approach is robust to different types of adversarial noise in HTML.
@@ 3 KnowPhish Construction
In this section, we introduce how to construct KnowPhish, a large-scale multimodal BKB. We start by conducting empirical analysis on phishing feeds from different periods to find prospective indicators to proactively search for potential phishing targets. With our empirical findings, we then describe how to automatically construct KnowPhish using a publicly-accessible multimodal knowledge graph.
@@ 3.1 Empirical Motivation
Question 1. How much do phishing targets change over time?
Question 2. What are the enduring characteristics shared by phishing feeds across different time intervals?
@@ 3.1.1 Data
To achieve this objective, we conducted a study using two distinct phishing datasets. The first dataset, $D_{1}$, includes phishing webpages that were collected three years ago by [32]. It encompasses a total of 29,496 phishing instances targeting 283 different brands. The second dataset, $D_{2}$, comprises phishing samples obtained from APWG[6] at the end of 2022. This dataset contains a total of 5,167 phishing examples targeting 391 unique brands.
Furthermore, we manually categorized each of these phishing targets into one of the following distinct industries, namely: 1) financial, 2) online service, 3) telecommunication, 4) e-commerce, 5) social media, 6) postal service, 7) government, 8) web portal, 9) video game, and 10) gambling. For the brands that cannot be classified into any of the ten categories, we categorize it as 11) other businesses. For instance, Bank of America is categorized into the financial category, while KFC is classified as other businesses.
@@ 3.1.2 Analysis
Observation 1. Phishing targets significantly change after three years.
In particular, in Figure 2 we observe significant changes in phishing targets when compared to three years ago, with only 118 of the 391 brands in $D_{2}$ being present in $D_{1}$. Some previous phishing targets such as EMS and De Volksbank no longer exist, while new targets like Bitkub, and USPS, have emerged. This trend of changes aligns with the findings from a previous study by [34].
Figure 2: Phishing target brands change substantially over time. The venn diagram shows the disparities in the phishing targets from the two phishing datasets, with a few phishing target examples provided for illustration.
@@
Observation 2. Even though a large number of emerging phishing targets appear, the industries of those phishing targets do not significantly change.
However, an intriguing revelation from Figure 3 is that despite the shifting landscape of phishing targets, the industries targeted by these phishing attacks have remained mostly consistent. For example, De Volksbank (in $D_{1}$), Bitkub, and Banco Do Brazil (in $D_{2}$) are all banking organizations, even though they became phishing targets at different times. Almost all the phishing targets in $D_{2}$ can still be categorized in the ten industries. The majority of these targets continue to belong to the same industries that have historically been the focus for phishing attacks, such as financial, telecommunication, and postal service, among others, while only a few brands such as KFC, Hydroquebec cannot be covered by the ten industries. We provide more examples of the phishing targets belonging to different industries in Table 8 in Appendix A.
It is worth noting that this observation echoes with the six Principles of Influence [10] upon which social engineering relies, namely the significance of authority in achieving successful persuasion. In the context of social engineering, threat actors, including phishing attackers, tend to focus their efforts on authorized and higher-value entities to maximize their gains from unlawful acquisition of sensitive information, rather than impersonate less-reputable and lower-value firms.

### v2

@@ Abstract
Phishing attacks have inflicted substantial losses on individuals and businesses alike, necessitating the development of robust and efficient automated phishing detection approaches. Reference-based phishing detectors (RBPDs), which compare the logos on a target webpage to a known set of logos, have emerged as the state-of-the-art approach. However, a major limitation of existing RBPDs is that they rely on a manually constructed brand knowledge base, making it infeasible to scale to a large number of brands, which results in false negative errors due to the insufficient brand coverage of the knowledge base. To address this issue, we propose an automated knowledge collection pipeline, using which we collect a large-scale multimodal brand knowledge base, KnowPhish, containing 20k brands with rich information about each brand. KnowPhish can be used to boost the performance of existing RBPDs in a plug-and-play manner. A second limitation of existing RBPDs is that they solely rely on the image modality, ignoring useful textual information present in the webpage HTML. To utilize this textual information, we propose a Large Language Model (LLM)-based approach to extract brand information of webpages from text. Our resulting multimodal phishing detection approach, KnowPhish Detector (KPD), can detect phishing webpages with or without logos. We evaluate KnowPhish and KPD on a manually validated dataset, and a field study under Singapore’s local context, showing substantial improvements in effectiveness and efficiency compared to state-of-the-art baselines.
@@ 1 Introduction
Phishing attacks are one of the most impactful types of scams, harming both individuals and businesses: in 2023, an estimated \$1.026 trillion was lost by consumers worldwide in scams [5], of which phishing scams are among the most common types. Phishing scams also account for about 90% of data breaches for organizations [12]. This problem has been exacerbated by the proliferation of automated phishing kits, enabling malicious actors to easily mimic genuine pages while evading detection using countermeasures [63]. Accordingly, the number of phishing attacks has increased by 47.2% in 2022 compared to the previous year [64]. These underscore the importance and urgency of tackling the issue, and the need for effective automatic phishing detection approaches.
@@
Many efforts have been made to counter phishing attacks, including anti-phishing blacklists, heuristic-based, and feature-based models [15]. Blacklist-based methods [42, 44, 47] compare an input URL with a predefined blacklist of malicious URLs, but these are reactive approaches that require phishing sites to be first reported or detected through other means. Heuristic-based [14] and feature-based models [29, 39, 19, 61, 32, 56, 37] extract features to proactively identify new phishing webpages. However, as these do not utilize logo information, they are greatly limited in their ability to detect plentiful phishing pages which are mainly identifiable by the presence of a logo [34]. Moreover, as they depend on statistical features for detecting phishing pages, they are susceptible to distribution shifts due to the constantly changing nature of phishing campaigns.
@@
In contrast, reference-based phishing detectors (RBPDs), which work by comparing the logos on a target webpage to a known set of logos, have been established as the state-of-the-art phishing detection paradigm, garnering considerable research attention [1, 33, 34]. Specifically, an RBPD consists of brand knowledge base (BKB) containing brand information (the logos and legitimate domains of brands) and a detector backbone which uses information from this BKB for phishing detection. To detect if a webpage is a phishing or benign page, RBPDs first identify the webpage’s brand intention, i.e., the brand that the webpage presents itself as (e.g., a webpage with an Adobe logo and appearance has the brand intention of Adobe). Then, if the webpage is detected to have an intention of a certain brand, but its domain does not match the legitimate domains of that brand, the webpage is classified as phishing. As virtually all phishing pages do not utilize the legitimate domains of the original brand, this approach typically obtains high precision. Moreover, as this approach is based on an invariant that does not change over time, it is relatively robust to distribution shift [33, 34].
@@
Challenges.  Despite the advantages of RBPDs, they have two major limitations which we focus on. (1) The first limitation is Limited-Scale BKB: RBPDs fundamentally rely on their BKB to identify the brand intention of a website. However, it is labor-intensive to construct and maintain a large-scale BKB manually. Phishpedia [33] and PhishIntention [34] rely on manual curation, hence are limited to a small BKB of 277 brands. A recent method, DynaPhish [35], proposes to dynamically expand the BKB during deployment time. However, this leads to extremely long running time, e.g., averaging $10.6$ seconds per sample in our experiments. Moreover, it may fail to construct brand knowledge of novel phishing targets if phishing pages’ logos are different from those displayed on legitimate pages. (2) The second limitation is Textual Brand Intention: Phishing webpages can convey their brand intention via text in HTML, instead of via logos. Existing RBPDs cannot identify it because they solely operate within the image modality.
@@
Present Work.  In this paper, we seek to address both of these issues in the context of a static environment, where input webpages are not interactable. First, we propose an automated knowledge collection pipeline, with which we construct a large-scale multimodal BKB named KnowPhish. KnowPhish is constructed based on our empirical analysis which finds that phishing targets mostly belong to a few high-value industries. Hence, using brand-industry relations modeled from a publicly available knowledge base, Wikidata[57], we search for a set of potential phishing targets and their brand knowledge predictively, leading to a BKB covering around 20k potential phishing targets. We also incorporate extra data sources such as Tranco domain list[46] and Google Image Search[18] for brand knowledge enhancement. Therefore, KnowPhish is equipped with rich logo, alias, and domain variants of each brand, which can be used to boost the performance of existing RBPDs in a plug-and-play manner.
@@
Next, to address the issue of textual brand intention, we develop a Large Language Model (LLM)-based approach to identify the brand intention of webpages in conjunction with the alias information in our KnowPhish BKB. Our approach can be directly integrated with any standard visual RBPD [34], augmenting it to form a multimodal RBPD which detects brand intention through both visual and textual modalities. Our resulting multimodal phishing detection approach, named KnowPhish Detector (KPD), can operate with static webpage data to detect phishing webpages with or without logos.
@@
We then evaluate the effectiveness and efficiency of KnowPhish and KPD on TR-OP, a manually validated dataset that comprises 5k benign and 5k phishing webpages. We also evaluate our approach on a field study under Singapore’s local context, to study how well different approaches generalize to such a local context. The resulting data, which we call SG-SCAN, contains 10k webpages from Singapore local webpage traffic over 6 months. In experiments on the two datasets, KnowPhish significantly boosts the effectiveness of various RBPDs, and is 30 or more times faster than the on-deployment framework DynaPhish [35], when equipped with image-based RBPDs. Moreover, incorporating our multimodal approach, KPD, can substantially boost the number of detected phishing webpages.
@@
  • Multimodal Brand Knowledge Base. We propose KnowPhish, a large-scale multimodal BKB for phishing detection, and its automated construction approach. KnowPhish can be used in any RBPDs to boost their brand knowledge in a plug-and-play manner.
  • Multimodal Reference-based Phishing Detector. We propose an LLM-based approach to identify textual brands from HTML to handle logo-less phishing webpages. Our approach directly integrates with any existing RBPD, augmenting it to form a multimodal RBPD, named KnowPhish Detector (KPD), which can detect phishing webpages with or without logos.
  • Effectiveness and Efficiency Extensive experiments show that KnowPhish significantly enhances the performance of various RBPDs, including KPD, while achieving much better runtime efficiency than DynaPhish. We also demonstrate their effectiveness in a field study and validate the robustness of KPD to adversarial attacks.
2 Motivating Examples
The development of KnowPhish and KPD is motivated by the brand knowledge construction difficulty of DynaPhish. Below, we present a brief workflow of DynaPhish, a few examples of where DynaPhish fails to build brand knowledge, and how KnowPhish and KPD address these difficulties.
DynaPhish expands brand knowledge through a two-step process when encountering a webpage with an unseen logo. (1) It first utilizes Google Search (GS) to validate the webpage’s domain popularity to determine its benignity. The logo-domain pair will be added to its brand knowledge if such validation passes. (2.1) If the validation fails, it recognizes the brand name of the logo using Google Logo Detector (GLD) and searches for the legitimate domain associated with that logo using another GS with the brand name as the query. (2.2) It then compares the input logo with the logo on each webpage from the GS results. If a match is found, the corresponding domain and logo are added to the brand knowledge.
However, DynaPhish may fail to construct new brand knowledge and correctly classify phishing pages due to the failure in step (2.2), i.e., logo-matching. A few examples of logos from phishing pages and the logos from their corresponding GS result pages are shown in Figure 1. The difference in appearance between the logos on phishing pages and GS result pages causes failure in the logo-matching process, even if GLD identifies the correct brand names, further resulting in brand knowledge construction failure and false negatives.
Figure 1: Comparison of the workflow between DynaPhish and KnowPhish to identify the brand intention of the two phishing page examples.
These examples motivate us to address such false negative problems in two ways. Firstly, we build KnowPhish from a high-quality data source, Wikidata, based on brand profiles (e.g., names and domains). This approach bypasses the logo-matching process used by DynaPhish and directly incorporates enhanced brand knowledge, such as logos and aliases. Therefore, the enriched logo variants increase the chance for image-based RBPDs to match the correct brand. For instance, the Swiss Post phishing page logo in Figure 1 is covered in KnowPhish, thus enabling the logo-matcher to identify its brand intention. Secondly, even if image-based RBPDs fail to match the logo, we remedy this by also using the text-brand extractor from KPD to identify the text brand intention from HTML and URL, in conjunction with the rich aliases in KnowPhish. As the Etisalat phishing page in Figure 1 shows, an LLM takes both the HTML and URL as input and outputs a predicted brand for the page. The predicted brand matches an alias from KnowPhish, thus the text-brand extractor detects its brand intention as the brand of that alias. We will elaborate on each in Sections 4 and 5, respectively.
@@ 3 Formalization
Threat Model.  In a phishing attack, a malicious actor misleads visitors into believing that their webpage comes from a legitimate brand, misleading users into providing their credentials (such as username and password). Note that we do not consider other types of attacks, such as malware or miscellaneous scams that do not fit the above description, to be within our scope.
@@
Reference-Based Phishing Detection.  Since phishing webpages need to convey brand intention, the state-of-the-art RBPD approach relies on identifying this brand intention, by comparing images on the page to a set of known reference logos. Formally, an RBPD consists of a brand knowledge base (BKB) and a detector backbone utilizing this BKB.
@@
$f_{\mathcal{B}}(w)=\begin{cases}\text{Phishing}&\text{if }g_{\mathcal{B}}(w)=b^{\prime}\text{ and }b^{\prime}\neq\varnothing\text{ and }\\ &\ \ \ w.\textit{domain}\notin b^{\prime}.\textit{domains}\\ \text{Benign}&\text{otherwise}\end{cases}$
Evasion Attacks.  The attackers may attempt to bypass $f_{\mathcal{B}}$ via the following methods:
T1: Phishing with Logo Variants. To circumvent the online knowledge expansion approach for $\mathcal{B}$ (e.g., DynaPhish[35]), attackers can use other legitimate logo variants of $b$ instead of the ones displayed on $b$’s official webpages.
T2: Phishing with Text Brands. Instead of using a logo to present its brand intention, the attacker can rely on text w.html to show its brand intention, making image-based phishing detectors completely fail.
T3: HTML-oriented Attacks. Phishing attackers may employ evasion techniques on HTML, such as typosquatting, prompt injection, and text-to-image attacks to hinder effective information extraction by text-based methods.
We address T1 by constructing a large-scale multimodal BKB with rich logo information, i.e., multiple logos per brand. We address T2 by developing an LLM-based approach to extract text brands from HTML. For T3, we show in Section 6.5 that our LLM-based approach is generally robust to different types of adversarial noises in HTML.
@@ 4 KnowPhish Construction
In this section, we introduce how to construct KnowPhish, a large-scale multimodal BKB. We start by conducting empirical analysis on phishing feeds from different sources and periods to find prospective indicators to search for potential phishing targets proactively. With our empirical findings, we then describe how to automatically construct KnowPhish using a publicly accessible multimodal knowledge graph.
@@ 4.1 Empirical Motivation
Question 1. Do phishing targets differ across different phishing feeds?
Question 2. What are the enduring characteristics shared by phishing feeds across different sources and periods?
@@ 4.1.1 Data
To achieve this objective, we conducted a study using two distinct phishing datasets. The first dataset, $D_{1}$, includes phishing webpages that were collected by [33] from OpenPhish[42] three years ago. It encompasses a total of 29,496 phishing instances targeting 283 different brands. The second dataset, $D_{2}$, comprises phishing samples obtained from APWG[7] at the end of 2022. This dataset contains a total of 5,167 phishing examples targeting 391 unique brands.
Furthermore, we manually categorized each of these phishing targets into one of the following distinct industries, namely: 1) financial, 2) online service, 3) telecommunication, 4) e-commerce, 5) social media, 6) postal service, 7) government, 8) web portal, 9) video game, and 10) gambling. For the brands that cannot be classified into any of the ten categories, we categorize them as 11) other businesses. For instance, Bank of America is categorized into the financial category, while KFC is classified as other businesses.
@@ 4.1.2 Analysis
Figure 2: Phishing targets change substantially. The Venn diagram shows the disparities in the phishing targets from the two phishing datasets, with a few phishing target examples provided for illustration.
@@
Observation 1. Phishing targets may exhibit variation depending on the datasets involved.
In particular, in Figure 2 we observe significant phishing target disparities between $D_{1}$ and $D_{2}$, with only 118 of the 391 brands in $D_{2}$ being present in $D_{1}$. This can be attributed to many potential factors, including temporal shifts and data collection methodologies. On the temporal shifts side, emerging phishing targets (e.g., Bitkub) may supplant existing ones over time[35]. On the data collection methodology side, OpenPhish employs proprietary detectors for phishing URL collection, whereas APWG relies on human reports. Such divergent approaches to data gathering directly impact the composition of phishing targets. These factors highlight the non-trivial challenge of manually managing the dynamic nature of phishing targets.
Observation 2. Despite the dynamic nature of phishing targets, the industries of those phishing targets remain mostly consistent.
However, an intriguing revelation from Figure 3 is that despite the shifting landscape of phishing targets, the industries targeted by these phishing attacks have remained mostly consistent. For example, De Volksbank (in $D_{1}$), Bitkub, and Banco Do Brazil (in $D_{2}$) are all banking organizations, even though they are phishing targets from different sources. Almost all the phishing targets in $D_{2}$ can still be categorized in the ten industries. The majority of these targets continue to belong to the same industries that have historically been the focus for phishing attacks, such as financial, telecommunication, and postal service, among others, while only a few brands such as KFC, Hydroquebec cannot be covered by the ten industries. Our supplementary materials[26] provide more examples of the phishing targets of different industries.
It is worth noting that this observation echoes the six Principles of Influence [11] upon which social engineering relies, namely the significance of authority in achieving successful persuasion. In the context of social engineering, threat actors, including phishing attackers, tend to focus their efforts on authorized and higher-value entities to maximize their gains from unlawful acquisition of sensitive information, rather than impersonate less-reputable and lower-value firms.